OpenStudy (anonymous):

Solve the following equation: x^2 − 5x + 6 = 0

OpenStudy (anonymous):

(x-3)(x-2)=0 (factor rule ) x-3= 0 =3 x-2=0 =2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np